OpenStudy (anonymous):

find the equation of the tangent plane surface at the specific points z=y cos (x - y) (2,2,2)

OpenStudy (goformit100):

it's east

OpenStudy (goformit100):

diffrentiate the given function with respect to x

OpenStudy (goformit100):

then put the value of the diffrential coefficient in the question

OpenStudy (goformit100):

for example it can be found in the same way when you deal with the circle chapter.

OpenStudy (perl):

i dont think so goformit, -thumbs down-

OpenStudy (perl):

he wants an equation, - gulps-

terenzreignz (terenzreignz):

@Andresfon12 Gradients... you know how to get them? :D

OpenStudy (goformit100):

(Y-y)^2 = dy/dx = (X-x)^2 is the tangent equation

OpenStudy (perl):

ok lets find the equation of the tangent plane at a point for z= f(x,y) -calls wikipedia-

terenzreignz (terenzreignz):

@Andresfon12 Place all expressions on one side of the equation (it doesn't really matter which side) and the other side should be 0. Let's call that function g(x,y,z) g(x,y,z)=z-y cos(x - y) = 0 Right?

OpenStudy (perl):

goformit100, you in mit?

OpenStudy (anonymous):

so far i know that fx= -(1)sin (x-y) and i think is fy=1 cos (x-y)

OpenStudy (perl):

goformit, ask for your money back, you suck

terenzreignz (terenzreignz):

I think you better redo that... anyway... since you've started this way, I think it best to redefine our g :) g(x,y,z) = -z + y cos(x - y)

terenzreignz (terenzreignz):

Now please find the partial derivatives of g with respect to x, y, and z...

OpenStudy (anonymous):

z=-1+y cos (x-y) , x= y cos (x-y) and y = -sin(x-y)

OpenStudy (perl):

the equation of the tangent plane is fx(x0,y0)(x−x0)+fy(x0,y0)(y−y0)−(z−z0)=0

OpenStudy (perl):

so we find fx, and fy , the partials with respect to x and y

OpenStudy (perl):

here i found it on this website with an example http://www.math.hmc.edu/calculus/tutorials/tangentplanes/

terenzreignz (terenzreignz):

Remember that all you need to define a plane in 3d is a point and a normal vector. By taking the gradient (the vector formed by the partial derivatives of g) and evaluating it at the point (2,2), you'll get the normal vector at that point, and since you already have a point, you'll get your plane.

OpenStudy (anonymous):

i know that z-2= (x- ) + (y- )

OpenStudy (perl):

given f(x,y) = y cos (x - y) fx = y(-sin(x-y)) fy= 1*cos(x-y) + y * (-sin(x-y))

OpenStudy (anonymous):

z=y

OpenStudy (perl):

given f(x,y) = y cos (x - y) fx = y(-sin(x-y)) = - y sin (x-y) fy= 1*cos(x-y) + y * (-sin(x-y)) = cos(x-y) - y sin(x-y) The equation will be fx (2,2) (x-2) + fy( y - 2) + (z-2) = 0

OpenStudy (anonymous):

i knew it product to find fy

OpenStudy (perl):

can you check my fx, and fy?

OpenStudy (anonymous):

product rule*

OpenStudy (perl):

yes

OpenStudy (perl):

yes he is right, I get z = 2

OpenStudy (perl):

because fx(2,2) = 0, and fy(2,2) = 0

OpenStudy (anonymous):

- y sin (x-y) =0 (2,2) cos(x-y) - y sin(x-y)= 1 (2,2)

OpenStudy (perl):

oh wait, sorry

OpenStudy (anonymous):

cos (0)=1

OpenStudy (perl):

fy(2,2) = 1 , so I get (1) (y-2) + (z-2) = 0

OpenStudy (perl):

the equation of tangent plane is fx (2,2)* (x-2) + fy(2,2) *( y - 2) + (z-2) = 0

OpenStudy (perl):

0(x-2) + 1(y-2) + z-2 = 0

OpenStudy (anonymous):

z-2=1(y-2)+(x-2)=> z-2=y=> z=y+2

OpenStudy (perl):

sorry made a mistake

OpenStudy (anonymous):

i notice

OpenStudy (perl):

the equation of tangent plane is fx (2,2)* (x-2) + fy(2,2) *( y - 2) - (z-2) = 0

OpenStudy (perl):

0 * (x-2) + 1( y-2 ) - (z-2) = 0

OpenStudy (anonymous):

z-2=y-2=> z=y

OpenStudy (perl):

correct

OpenStudy (perl):

:)

OpenStudy (anonymous):

thx @perl